OpenStudy (anonymous):

is this equation 7x+6-4x+12+3x-8 always,sometimes, or never true?

OpenStudy (anonymous):

x= -50

OpenStudy (anonymous):

I need the work for it.

OpenStudy (anonymous):

think bout it wat it askin sometimes

OpenStudy (anonymous):

wait huh?

OpenStudy (anonymous):

sometimes

OpenStudy (anonymous):

whats the combine like terms part?

OpenStudy (radar):

An equation requires an equal sign.

OpenStudy (anonymous):

there is a equal sign.

OpenStudy (radar):

Where is the equal sign?

OpenStudy (anonymous):

the question says.... is the equation 7x+6-4x=12+3x-8 always,sometimes, or never true? combine like terms on each side of the equation +6=3x+ now solve the equation.

OpenStudy (anonymous):

thats word for word of the question^

OpenStudy (radar):

O.K. I see the equal sign now, it wasn't in the original posted problem. 3x + 6 = 3x + 4 Does that look like it ever could be true??? Don't think so.

OpenStudy (anonymous):

so what do we put next?

OpenStudy (radar):

It is never true, subtract 3x from both sides you get 6=4 Not in Missouri

OpenStudy (anonymous):

i thought it was sometimes?

OpenStudy (radar):

Get those kind of thoughts out of your head...

OpenStudy (anonymous):

okay. how do i combine like terms with +6+3x+ ?

OpenStudy (anonymous):

@radar @Lyssa123

OpenStudy (anonymous):

wat u need me to do ???

OpenStudy (anonymous):

7x+6-4x=12+3x-8 always,sometimes, or never true? <-- first question combine like terms on each side of equation. +6=3x+ <---- second question the equation is always,sometimes, or never true <---- third question @Lyssa123 @radar

OpenStudy (radar):

Combine terms, means only "similar" terms. You can't combine 7x and 6. 7x + 6 -4x = 12 +3x - 8 combining similar terms on lefts gets 3x + 6 = 12 + 3x - 8, now combining similar terms on the right side get 3x + 6 = 3x + 4 Subtract 3x from both sides to isolate x 3x-3x + 6 = 3x - 3x + 4 0 + 6 = 0 + 4 6 = 4 That tell you it is a bogus equation and there is no solution.

OpenStudy (anonymous):

so is that never true or sometimes?

OpenStudy (radar):

I can't think of anytime when 4 equals 6, so I'm gonna say "never true"